OpenStudy (anonymous):

medal/fan please help Which system of equations can be used to solve the following problem? Each child ticket for a ride costs $3, while each adult ticket costs $5. If the ride collected a total of $150, and 40 tickets were sold, how many of each type of ticket were sold? Let c be the number of child tickets and a be the number of adult tickets.

OpenStudy (anonymous):

a

OpenStudy (anonymous):

b

OpenStudy (anonymous):

c

OpenStudy (anonymous):

d

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

can you help me with some more @princeharryyy

OpenStudy (princeharryyy):

D it is , @Soniatonia

OpenStudy (princeharryyy):

ok

OpenStudy (anonymous):

,Which equation shows the substitution method being used to solve the system of linear equations? x + y = 6 x = y + 5 A. x = (x – 6) + 5 B. (y + 5) + y = 6 C. x + (y + 5) = 6 D. x + y = y + 5

OpenStudy (princeharryyy):

B @Soniatonia

OpenStudy (anonymous):

The system of equations is coincident. What are the missing values?

OpenStudy (anonymous):

8x + y

OpenStudy (princeharryyy):

8x + 10y = 16

OpenStudy (anonymous):

What is the best approximation of the solution to the system, to the nearest integer values?

OpenStudy (princeharryyy):

(-1.8,2.8)

OpenStudy (princeharryyy):

@Soniatonia x=-1.8, y=2.8

OpenStudy (anonymous):

How many solutions are there to the following system of equations? 3x – 9y = 0 –x + 3y = –3 A. infinitely many B. 2 C. 0 D. 1

OpenStudy (princeharryyy):

C 0

OpenStudy (princeharryyy):

@Soniatonia

OpenStudy (anonymous):

last one

OpenStudy (anonymous):

What is the solution to the system of equations?

OpenStudy (princeharryyy):

x=-2. y=1 @Soniatonia

OpenStudy (anonymous):

thank you sooo much :)

OpenStudy (princeharryyy):

welcome @Soniatonia